A NEW ONE-ELECTRON EFFECTIVE POTENTIAL FOR CaF BASED ON AB-INITIO CALCULATIONS

\underline{STEPHEN L. COY}, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; BRYAN M. WONG, Sandia National Laboratory, Livermore, CA 94551-0969; and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

The Rydberg spectrum of CaF has been extensively analyzed both theoretically and experimentally, and new QDT fit and Stark/polarization results will soon be reported. In these results, a new level of detail is achieved that enables a comparison between experimental results and reaction matrix calculations for state decompositions, and dependences on R(Ca-F), as well as energies. The most commonly used potential for the Rydberg states of CaF was presented a decade ago^{a}. It has been effective in predicting energies for electronic states with R(Ca-F) near the equilibrium separation, but does not include any variation with internuclear separation, and predicts quantities such as the ion-core dipole moment that disagree with ab-initio results as a function of R. We have performed all-electron ccsd(T) calculations for CaF^{+} and test-charge potential calculations for a wide range of R(Ca-F) values and test-charge angular positions. We will discuss the differences between these new results and the earlier potential and describe a new one-electron effective potential for CaF.

aM. Arif, Ch. Jungen and A. L. Roche J. Chem. Phys. \underline{106}(10), 4102 (1997).